A moving charge or current produces what kind of f All parts needed
Answer #1

a) magnetic field

b) the charged particle will remain at rest.

because, there is no magnetic force on ststaionary charged particle.

c) 4.5*10^-5 T

d) zero.

because, they produce equal magnetic field but opposite in direction. so, net magnetic field will be zero.

e)

i) no, because magnetic force always perpendicular to instantaneous dispalcement.

W = F*d*cos(90) = 0

ii) uniform circular motion.

f) The direction of induced emf(or current) is always such as to oppose the change in magnetic flux that genereted it.

induced emf = -A*dB/dt

the negative sign indicates Lenz'slaw.